OpenStudy (anonymous):

HEY THERE PLS HELP ME: Integers as Exponents 1. explain why (-1)^n=1 for any even # 2. Explain why both 4x^0-3y^0 & (4x-3y)^0 simpifyy to 1

OpenStudy (neer2890):

\[(-1)^{n}=1\] for any n=even because power means nothing but multiplication. for e.g. n=2 (-1)^2=(-1)*(-1)=1 similarly you can check for any other even value of n.

OpenStudy (neer2890):

Also, anything raise to even power is always positive. and 1 raise to power anything is always 1 itself.

OpenStudy (neer2890):

Anything raise to power zero is always 1 because Let us assume without the loss of generality, a,m,n belongs to R(real numbers) and a not equal to 0. then \[\frac{ a ^{m} }{ a ^{n} }=a ^{m-n}\] when m=n \[\frac{ a ^{n} }{ a ^{n} }=a ^{n-n}\] a^n cancels out on left side and right side becomes a^0 and it implies \[1=a ^{0}\] hence proved.

OpenStudy (anonymous):

wait what do u mean by the left side and the right side